OpenStudy (anonymous):

you can drive your car 2.17 miles with one gallon of gasoline. At that rate, how many miles can you drive with a3.2 gallons of gasoline?

OpenStudy (anonymous):

6.944 miles in 3.2gallons

OpenStudy (texaschic101):

2.17/1 = x/3.2 (2.17 miles on 1 gal = x miles on 3.2 gall) cross multiply (1)(x) = (3.2)(2.17) x = 6.944 miles if you have to round to the nearest whole number, answer is 7 miles

OpenStudy (anonymous):

thank you!

OpenStudy (texaschic101):

anytime :)